# Out-of-Distribution Detection with a Single Unconditional Diffusion Model

### v1

@@ 4 Related Works
Modern OOD detection can be divided roughly into three categories: feature-based, likelihood-based, reconstruction-based. Feature-based methods extract features from ID samples and fit a likelihood or distance function as an OOD detector. For instance, one can obtain the latent representations of a test sample using an autoencoder and measure its Mahalanobis distance to the representations of ID samples [24]. Distances between features derived from self-supervised learning models are also utilized in similar contexts [25, 26]. Similar to our work, Xiao et al. [10] showed that one can perform OOD detection using features from a single discriminative model trained on ImageNet.
Likelihood-based approaches leverage a generative model trained on ID samples. These methods typically employ variants of the log-likelihood of a test sample under the model as the OOD detection score. Nalisnick et al. [2] first pointed out that deep generative models might erroneously assign higher likelihoods to OOD samples. Several explanations have been proposed, such as the input complexity [16] and typicality [27] of samples. As a result, just as we show in Sec. 3.1, vanilla likelihoods are rarely used. Instead, variants derived from the log-likelihood have been proposed, such as likelihood ratios [4], ensembles of the likelihood [5], density of states [6], energy-based models [28] and typicality tests [27]. A similar line of work to ours is MSMA [29], which uses the score function over discrete noise levels for OOD detection. One can view MSMA as a specific case of DiffPath, where the former only utilizes the first-order statistic, while we generalize to higher-order terms. MSMA proposes to measure the score at various noise levels, while our method sums over the entire diffusion path. It is worth emphasizing that MSMA requires different models for different ID tasks, unlike our single model setup.
Reconstruction-based approaches evaluate how well a generative model, trained on ID data, can reconstruct a test sample. Earlier works utilize autoencoders [30] and GANs [31] for reconstruction. Recent works have adapted unconditional DMs to this approach due to its impressive sample quality. A test sample is first artificially corrupted before being reconstructed using the DM’s sampling process. DDPM-OOD [7] noises a sample using the forward process and evaluates the perceptual quality [32] of the reconstructed sample. Projection Regret [9] adopts a similar approach, but uses a Consistency Model [33] and introduces an additional projection regret score. LMD [8] corrupts the image by masking and reconstructs the sample via inpainting. Evidently, DiffPath differs from these diffusion approaches as we do not utilize reconstructions. We also stress again that these baselines require different models for different ID tasks.
@@ 5 Experiments
Table 3: AUROC scores for various in-distribution vs out-of-distribution tasks. Higher is better. Bold and underline denotes the best and second best result respectively. We also show the number of function evaluations (NFE) for diffusion methods, where lower is better.
@@
DiffPath-1D 0.965 0.998 0.551 0.966 0.971 1.00 0.972 0.992 0.998 1.00 0.994 0.827 0.936 50
DiffPath-1D 0.966 0.999 0.552 0.975 0.971 1.00 0.973 0.988 0.999 1.00 0.998 0.851 0.939 100
DiffPath-6D 0.936 0.997 0.596 0.997 0.921 0.999 0.932 0.999 0.997 1.00 0.995 0.935 0.942 50
DiffPath-6D 0.920 0.999 0.593 0.999 0.924 0.999 0.936 1.00 0.999 1.00 0.998 0.946 0.943 100
@@
We evaluate pairwise OOD detection using CIFAR10 (C10), SVHN, and CelebA as ID datasets, and CIFAR100 (C100) and Textures as additional OOD datasets. We employ an unconditional diffusion model trained on ImageNet at $64\times 64$ resolution, using the Improved DDPM’s [34] open-source checkpoint [^2] https://github.com/openai/improved-diffusion. All datasets are resized to $64\times 64$ when evaluating with DiffPath (see Sec. 5.2 for ablations on resizing techniques).
@@
Table 4: Ablation results when we vary $p_{0}({\mathbf{x}})$, the distribution the single DM is trained on. We use DiffPath-1D with 50 NFEs. Bold and underline denotes the best and second best result.
@@
$p_{0}({\mathbf{x}})$ SVHN CelebA C100 Textures C10 CelebA C100 Textures C10 SVHN C100 Textures Average
C10 0.862 0.359 0.538 0.724 0.788 0.807 0.803 0.862 0.658 0.927 0.704 0.839 0.739
SVHN 0.975 0.394 0.512 0.698 0.985 0.997 0.984 0.960 0.625 0.996 0.621 0.779 0.794
CelebA 0.915 0.55 0.519 0.693 0.881 0.779 0.872 0.840 0.772 0.902 0.789 0.850 0.780
ImageNet 0.965 0.998 0.551 0.966 0.971 1.00 0.972 0.992 0.998 1.00 0.994 0.827 0.936
@@ 5.1 Main Results
Table 3 summarizes our main results. Both variants of DiffPath, with 50 and 100 NFEs, obtain the top four best average results across all 12 tasks, with DiffPath-6D at 100 NFEs performing the best. Interestingly, dropping to 50 NFEs only reduces performance marginally while improving inference speed by 2×2$\times$2 ×. Specifically, DiffPath-6D at 100 NFEs achieves the best result in six out of twelve tasks, and second best result on one of them, while at 50 NFEs achieves second best result in four out of twelve tasks. Despite using only a scalar statistic for OOD detection, DiffPath-1D at 100 NFEs still obtains the best result in five out of twelve tasks. Thus, our results validate the efficiency and versatility of DiffPath on diverse OOD tasks.
All variants of DiffPath struggle with the C10 vs C100 task due to similarities between these datasets, scoring below 0.7 AUROC. However, this task is also challenging for all baselines, with only Improved CD and IC achieving AUROC scores above 0.7. Interestingly, DiffPath’s performance is strongest when either ID or OOD dataset is CelebA. For instance, we achieve perfect score when distinguishing against SVHN across both variants, and near perfect score (greater than $0.99$) against C10/C100. This corroborates the result in Fig. 2, which shows that the histogram of CelebA is well separated from the other datasets.
Interestingly, DiffPath maintains consistent performance when swapping ID and OOD datasets. As a single model is used, the statistics extracted for a dataset is identical (up to differences between train test split) regardless of whether it is ID or OOD. Other methods do not exhibit such behavior as they require different generative models for different tasks. For example, in the C10 vs CelebA/CelebA vs C10 tasks, DiffPath-1D achieves identical scores of 0.999/0.999, while MSMA and LMD achieves scores of 0.675/0.871 and 0.557/0.989 respectively. We believe this symmetry is desirable as it is similar to binary classification, where performance is unchanged when labels are swapped.
DiffPath requires fewer NFEs per sample than most diffusion models, making it faster during inference. Only MSMA uses fewer NFEs, but it performs worse and needs different models for each ID task. Reconstruction-based methods like DDPM-OOD and LMD require many reconstructions per sample, leading to high NFE requirements. In particular, LMD requires a hefty $10^{4}$ NFEs as the original implementation requires 10 reconstructions per sample, with each utilizing 1000 NFEs. As for NLL and IC, both underperform our method despite our experiments showing that roughly 1000 NFEs are required to obtain an accurate estimate of the likelihood.
Figure 4: Plots of AUROC performance for various ablation experiments. Average performance over all tasks when the number of DDIM steps is varied (left) and when the image resizing technique is varied (right). We use DiffPath-6D with 50 NFEs for both. Full results in Sec. B.1 of the appendix.
@@ 5.2 Ablations
Here we quantitatively justify our claim in Sec. 3.2 that the data distribution used to train the DM, $p_{0}({\mathbf{x}})$, has to be large and diverse for OOD detection. We report ablation results in Table. 4 using DMs trained on CIFAR10, SVHN, CelebA and ImageNet. The results show that the DM trained on ImageNet significantly outperforms the rest, achieving the best result across 8 out of 12 tasks, second best on 3 tasks and the best average result by a large margin. This is despite the other variants being trained on samples seen in the experiments, while ImageNet was not. This suggests that it is unnecessary for generative models to be trained on in-distribution data for OOD detection; rather, the model should be trained to learn diverse and general representations.
@@
We investigate how the performance of our method varies with the number of DDIM steps used in Fig. 4. Interestingly, the best results are obtained at 25, 50 and 100 DDIM steps, with near identical performance. As the experiments with 50 and 100 steps are more consistent throughout tasks (see Sec. B.1 in appendix for full ablation results), we report results for 50 and 100 NFEs in our main experiments in Table. 3. We hypothesize that the slight decrease in performance at larger DDIM steps is due to the use of the finite difference in Eq. 10, where the difference in consecutive predictions $\bm{\epsilon}(\cdot,t+\Delta t)-\bm{\epsilon}(\cdot,t)$ shrinks rapidly. We leave the investigation of this phenomena in greater detail to future work.
Image Resizing Technique.
Since we utilize a single DM trained on ImageNet at $64\times 64$ resolution, samples from image distributions of different resolutions have to be resized appropriately. As such, we ablate for resizing techniques that are implemented in the PyTorch library. The results in Fig. 4 show that DiffPath is relatively robust, with average AUROC approximately 0.9 or higher across all resizing techniques considered. However, bicubic and bilinear interpolation outperform the rest, with the latter obtaining the best performance. Thus, we use bilinear interpolation for the experiments described in this work.
@@ 6 Conclusion
In this work, we proposed Diffusion Paths (DiffPath), a method of OOD detection using a single diffusion model by characterizing properties of the forward diffusion path. In light of the growing popularity of foundational generative models, our work demonstrates that a generalist model–one that has learnt diverse representations from data–can also be applied to out-of-distribution detection. There are several interesting future directions that arise from our work; for instance, applying DiffPath to other modalities, given that diffusion modelling has been extended to domains such as video, audio, language and time series, as well as investigating if higher-order terms in the Taylor expansion, or leveraging different instantiations of the PF ODE might lead to better performance. We hope that our work sparks greater interest among practitioners in utilizing other foundational generative models beyond DMs for OOD detection, such as large language models.
@@
We only calculate the simple time derivative and found that it works well experimentally, although one might compute the full derivative to quantify the curvature completely. We leave this for future work. Also, we consider a DM trained on ImageNet, which may not be general enough for specialized applications such as medical images. For such purposes, one could consider including domain-specific data when training the DM, or utilize larger DMs such as Stable Diffusion.
@@ 7 Acknowledgements
We would like to thank Pranav Goyal for help with the experiments. This research is supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-RP-2020-017).
Supplementary Material for “Out-of-Distribution Detection with a Single Unconditional Diffusion Model"
@@ A.1 Theorem 1
This assumption essentially states that both marginals have at most polynomial growth at infinity and thus can be normalized.
@@ Appendix B Experimental Details
As mentioned in Sec. 5, we utilize a single unconditional diffusion model trained on ImageNet at $64\times 64$ resolution. The checkpoint that we use is trained using Improved DDPM’s $L_{\textrm{hybrid}}$ objective (Eq. 16 of Nichol and Dhariwal [34]). The model uses a cosine noise schedule with a total of 4000 diffusion steps. The checkpoint is publicly available on the official GitHub repository [^3] https://github.com/openai/improved-diffusion. For DiffPath-1D, we fit a KDE using a Gaussian kernel with a bandwith of 5. For DiffPath-6D, we fit a GMM using 50 mixture components, each with diagonal covariance. Both are implemented using the sklearn library. We calculate the diffusion statistics using a single A5000 GPU.
Diffusion Baselines
@@
For all diffusion baselines, we rely on the official GitHub repositories and open-source checkpoints where possible. The repositories are listed as follows: MSMA [^4] https://github.com/ahsanMah/msma, DDPM-OOD [^5] https://github.com/marksgraham/ddpm-ood, LMD [^6] https://github.com/zhenzhel/lift_map_detect. For NLL and IC, we either use the pre-trained CIFAR10 checkpoint from Improved DDPM or train separate models for SVHN and CelebA using the same hyperparameters at $32\times 32$ resolution. We calculate the NLL using the default implementation in Improved DDPM, while we compute IC using code from the LR repository [^7] https://github.com/XavierXiao/Likelihood-Regret due to lack of official code from the IC authors. We train all baselines using 1-3 A5000 GPUs.
B.1 Full Results for Ablations
We provide full results of the ablations presented in Fig. 4 in Table 5 and Table 6. When varying the NFE in Table 5, we observe that certain tasks are more sensitive than others, although the average across all tasks is much more stable. For instance, the worst performance between CIFAR10 vs SVHN has AUROC of 0.834 at 500 steps, while the best performance has AUROC of 0.936 at 50 steps. However, other tasks remain unchanged when NFE is varied, such as SVHN vs Textures or CelebA vs SVHN. We generally find 50 and 100 steps to the sweet spot.
When the resizing method is varied in Table 6, we find bicubic and bilinear interpolation to perform significantly better than the rest, with bilinear achieving the best scores in most tasks. As such, we use bilinear in all other experiments in the main text.
Table 5: Full ablation results when we vary the number of DDIM steps (NFE). We use DiffPath-6D. Bold and underline denotes the best and second best result.
C10 vs SVHN vs CelebA vs
NFE SVHN CelebA C100 Textures C10 CelebA C100 Textures C10 SVHN C100 Textures Average
25 0.917 0.996 0.587 0.997 0.919 0.999 0.929 1 0.995 1 0.993 0.970 0.942
50 0.936 0.997 0.596 0.997 0.921 0.999 0.932 0.999 0.997 1 0.995 0.935 0.942
100 0.920 0.999 0.593 0.999 0.924 0.999 0.936 1 0.999 1 0.998 0.946 0.943
200 0.882 1 0.588 0.999 0.912 1 0.926 1 0.999 1 0.999 0.973 0.940
500 0.834 0.999 0.583 0.999 0.850 1 0.870 1 0.999 1 0.998 0.981 0.926
Table 6: Full ablation results when we vary the resizing method of the images. We use DiffPath-6D. Bold and underline denotes the best and second best result.
C10 vs SVHN vs CelebA vs
Resizing Method SVHN CelebA C100 Textures C10 CelebA C100 Textures C10 SVHN C100 Textures Average
Lanczos 0.819 0.997 0.593 0.994 0.891 0.996 0.910 0.996 0.997 1 0.996 0.937 0.927
Nearest 0.815 0.995 0.575 0.973 0.836 0.999 0.865 0.993 0.979 0.995 0.983 0.983 0.916
Bicubic 0.900 0.988 0.589 0.987 0.899 0.994 0.916 0.996 0.986 1 0.979 0.930 0.930
Hamming 0.823 0.898 0.575 0.937 0.852 0.990 0.877 0.985 0.835 0.986 0.862 0.938 0.880
Box 0.814 0.974 0.574 0.957 0.835 0.995 0.866 0.991 0.951 0.987 0.962 0.938 0.904
Bilinear 0.936 0.997 0.596 0.997 0.921 0.999 0.932 0.999 0.997 1 0.995 0.935 0.942

### v2

@@ 4 Related Works
Modern OOD detection can be divided roughly into three categories: feature-based, likelihood-based, reconstruction-based. Feature-based methods extract features from inlier samples and fit a likelihood or distance function as an OOD detector. For instance, one can obtain the latent representations of a test sample using an autoencoder and measure its Mahalanobis distance to the representations of inlier samples [24]. Distances between features derived from self-supervised learning models are also utilized in similar contexts [25, 26]. Similar to our work, Xiao et al. [10] showed that one can perform OOD detection using features from a single discriminative model.
Likelihood-based approaches leverage a generative model trained on inlier samples. These methods typically employ variants of the log-likelihood of a test sample under the model as the OOD detection score. Nalisnick et al. [2] first pointed out that deep generative models might erroneously assign higher likelihoods to outlier samples. Several explanations have been proposed, such as the input complexity [16] and typicality [27] of samples. As a result, just as we show in Sec. 3.1, vanilla likelihoods are rarely used. Instead, variants derived from the log-likelihood have been proposed, such as likelihood ratios [4], ensembles of the likelihood [5], density of states [6], energy-based models [28] and typicality tests [27]. Diffusion Time Estimation [29] estimates the distribution over the diffusion time of a noisy test sample and uses the mean or mode as the OOD score. MSMA [30] uses the score function over discrete noise levels for OOD detection. One can view MSMA as a specific case of DiffPath which only utilizes the first-order statistic, while we generalize to higher-order terms. MSMA proposes to measure the score at various noise levels, while our method sums over the entire diffusion path. It is worth emphasizing that MSMA requires different models for different inlier distributions, unlike our single model setup.
Reconstruction-based approaches evaluate how well a generative model, trained on in-distribution data, can reconstruct a test sample. Earlier works utilize autoencoders [31] and GANs [32] for reconstruction. Recent works have adapted unconditional DMs to this approach due to its impressive sample quality. A test sample is first artificially corrupted before being reconstructed using the DM’s sampling process. DDPM-OOD [7] noises a sample using the forward process and evaluates the perceptual quality [33] of the reconstructed sample. Projection Regret [9] adopts a similar approach, but uses a Consistency Model [34] and introduces an additional projection regret score. LMD [8] corrupts the image by masking and reconstructs the sample via inpainting. Evidently, DiffPath differs from these diffusion approaches as we do not utilize reconstructions. We also stress again that these baselines require different models for different inlier tasks.
@@ 5 Experiments
Based on our earlier analysis, we hypothesize that DiffPath can be utilized for OOD detection across diverse tasks using a single model. In this section, we validate our hypothesis with comprehensive experiments across numerous pairwise OOD detection tasks and compare DiffPath’s performance against state-of-the-art baselines.
Table 3: AUROC scores for various in-distribution vs out-of-distribution tasks. Higher is better. DiffPath-6D-ImageNet and DiffPath-6D-CelebA denote our method using diffusion models trained with ImageNet and CelebA as base distributions respectively. Bold and underline denotes the best and second best result respectively. We also show the number of function evaluations (NFE) for diffusion methods, where lower is better.
@@
DiffPath-6D-ImageNet 0.856 0.502 0.580 0.841 0.943 0.964 0.954 0.969 0.807 0.981 0.843 0.964 0.850 10
DiffPath-6D-CelebA 0.910 0.897 0.590 0.923 0.939 0.979 0.953 0.981 0.998 1.00 0.998 0.999 0.931 10
@@
All experiments are conducted as of pairwise OOD detection tasks using CIFAR10 (C10), SVHN, and CelebA as inlier datasets, and CIFAR100 (C100) and Textures as additional outlier datasets. Unconditional diffusion models are employed at resolutions of $32\times 32$ and $64\times 64$. The model utilizing ImageNet as the base distribution is trained at $64\times 64$ resolution, while all other models are trained at $32\times 32$.
@@
Table 4: Ablation results when we vary $p_{0}({\mathbf{x}})$, the distribution the single DM is trained on. We use DiffPath-6D with 10 NFEs. Random denotes a randomly initialized model.
@@
$q_{0}({\mathbf{x}})$ SVHN CelebA C100 Textures C10 CelebA C100 Textures C10 SVHN C100 Textures Average
C10 0.939 0.484 0.604 0.870 0.961 0.961 0.973 0.982 0.719 0.997 0.796 0.950 0.853
SVHN 0.742 0.482 0.579 0.872 0.991 0.994 0.992 0.989 0.706 0.974 0.769 0.961 0.838
CelebA 0.910 0.897 0.590 0.923 0.939 0.979 0.953 0.981 0.998 1.00 0.998 0.999 0.931
ImageNet 0.856 0.502 0.580 0.841 0.943 0.964 0.954 0.969 0.807 0.981 0.843 0.964 0.850
Random 0.338 0.426 0.538 0.31 0.665 0.592 0.693 0.471 0.577 0.411 0.612 0.381 0.501
@@ 5.1 Main Results
Table 3 summarizes our main results. Here, we report outcomes for DiffPath-6D using ImageNet and CelebA as base distributions. DiffPath-6D-CelebA achieves an average AUROC of 0.931, comparable to the leading diffusion-based approach MSMA and outperforming all other baselines, while utilizing only a single model. Similar to MSMA, we attain this result using 10 NFEs, significantly surpassing other diffusion baselines that require an order of magnitude or more NFEs. When using ImageNet as the base distribution, the average AUROC of 0.850 is competitive with LMD, which requires several magnitudes more NFEs due to multiple reconstructions. This is despite the ImageNet model never having seen any samples from the evaluated distributions during training.
The empirical results indicate that the CelebA-based model outperforms the ImageNet-based model primarily due to its superior performance on tasks involving CelebA samples, whether they are in-distribution or out-of-distribution. For instance, DiffPath-6D-CelebA achieves near-perfect performance on all tasks where CelebA is in-distribution (rightmost columns), and in the CIFAR10 vs CelebA task. On tasks that do not involve CelebA samples, the two models exhibit roughly comparable performance. This suggests that distinguishing CelebA from other samples is particularly challenging, and that DiffPath benefits from exposure to inlier samples from the respective distributions during training. Next, we discuss the effect of the base datasets and other design considerations via ablations.
Table 5: Ablation on the number of DDIM steps (NFE). We use DiffPath-6D-CelebA.
C10 vs SVHN vs CelebA vs
NFEs SVHN CelebA C100 Textures C10 CelebA C100 Textures C10 SVHN C100 Textures Average
5 0.916 0.928 0.584 0.900 0.955 0.940 0.960 0.973 0.999 1.00 0.998 0.997 0.929
10 0.910 0.897 0.590 0.923 0.939 0.979 0.953 0.981 0.998 1.00 0.998 0.999 0.931
25 0.898 0.866 0.578 0.933 0.882 0.997 0.906 0.979 0.996 1.00 0.995 0.996 0.919
50 0.896 0.831 0.575 0.931 0.853 0.996 0.879 0.974 0.991 1.00 0.991 0.997 0.910
Table 6: Ablation results comparing DiffPath-1D and DiffPath-6D using models trained with ImageNet and CelebA as base distributions.
C10 vs SVHN vs CelebA vs
Method SVHN CelebA C100 Textures C10 CelebA C100 Textures C10 SVHN C100 Textures Average
DiffPath-1D-ImageNet 0.965 0.394 0.551 0.685 0.971 0.986 0.972 0.949 0.693 0.991 0.721 0.797 0.806
DiffPath-6D-ImageNet 0.856 0.502 0.580 0.841 0.943 0.964 0.954 0.969 0.807 0.981 0.843 0.964 0.850
DiffPath-1D-CelebA 0.956 0.811 0.545 0.688 0.948 0.690 0.933 0.932 0.899 0.666 0.881 0.911 0.822
DiffPath-6D-CelebA 0.910 0.897 0.59 0.923 0.939 0.979 0.953 0.981 0.998 1.00 0.998 0.999 0.931
@@ 5.2 Ablations
We investigate the impact of the base dataset on the performance of DiffPath-6D. In Table 4, we compare four different base distributions alongside a randomly initialized model. As a baseline check, we observe that the average AUROC of the randomly initialized model is 0.501, which is consistent with random guessing. This indicates that training on a base distribution is essential for the model to learn features for effective OOD detection.
Our ablations include CIFAR10 and SVHN as base distributions, in addition to CelebA and ImageNet shown in Table 3. Among these four base distributions, CelebA yields the best performance overall. Notably, the models trained on SVHN and CelebA demonstrate superior results when the inlier data aligns with their respective training distributions. This supports the established principle of training models on in-distribution samples, and we show that DiffPath-6D similarly benefits from such training. However, we underscore the key finding of our work: while in-distribution training enhances performance, it is not strictly necessary. DiffPath-6D exhibits strong performance even on tasks involving samples from distributions that the model has not encountered during training.
DiffPath 1D vs 6D.
Here we ablate on the choice of DiffPath-1D and DiffPath-6D. Table 6 shows that DiffPath-6D performs better than its 1D counterpart for both choices of base distributions. We attribute this to the increased robustness of aggregating multiple statistics, c.f. Sec. 3.5, and recommend practitioners to use DiffPath-6D in general. However, DiffPath-1D outperforms DiffPath-6D in certain instances. For instance, for the ImageNet model, DiffPath-1D achieves the best performance on CIFAR10 vs SVHN and in three out of four tasks where SVHN is the inlier distribution.
@@
We investigate how the performance of our method varies with the number of NFEs (DDIM steps) in Table 5. Overall, the changes in average AUROC are relatively minor as the NFEs are varied, suggesting that DiffPath is robust to the number of integration steps. The best result is obtained at 10 NFEs. While the finite difference approximation of the derivative (Eq. 10) should become more accurate as the number of NFEs increases, the aggregation of multiple statistics involving scores and its derivatives makes this effect less pronounced. We leave the investigation of this phenomena in greater detail to future work.
Table 7: Difference in performance when the incorrect resizing technique is used, which leads to overly optimistic results. Results on DiffPath-6D with ImageNet $64\times 64$ as the base distribution. The results with asterisk (*) denote the scores that have been computed inaccurately.
C10 vs SVHN vs CelebA vs
Correct Resizing SVHN CelebA C100 Textures C10 CelebA C100 Textures C10 SVHN C100 Textures Average
No 0.856 0.999* 0.580 0.999* 0.943 1.00* 0.954 1.00* 0.998* 1.00* 0.998* 0.981 0.942
Yes 0.856 0.502 0.580 0.841 0.943 0.964 0.954 0.969 0.807 0.981 0.843 0.964 0.850
5.3 Proper Image Resizing with a Single Model
Using a single DM with a fixed input resolution necessitates resizing all images to match the model’s resolution during evaluation. However, when datasets have differing original resolutions, naive resizing—upsampling lower-resolution images and downsampling higher-resolution ones—can lead to evaluation inaccuracies. Specifically, upsampling introduces blurriness due to pixel interpolation, while downsampling does not. This discrepancy allows the model to differentiate samples based on image blur rather than semantic content, resulting in overly optimistic performance metrics.
For instance, when evaluating DiffPath-6D trained at $64\times 64$ pixels on the CelebA vs CIFAR10 task, CIFAR10 images are upsampled (introducing blur) while CelebA images are downsampled. This imbalance enables trivial distinction between the samples. As illustrated in the first row of Table 7, tasks where only one distribution undergoes upsampling yield artificially high AUROC scores.
To mitigate this issue, we propose equalizing the resizing process by first downsampling higher-resolution images to the lower resolution of the other distribution, then upsampling both to the model’s required resolution. In the CelebA vs CIFAR10 example, CelebA images are downsampled to $32\times 32$ pixels before both samples are upsampled to $64\times 64$ pixels. This method ensures consistent blurring effects across all samples, reducing confounding factors. The second row of Table 7 demonstrates more accurate evaluations using this approach. We adopt this resizing procedure in all relevant experiments to ensure fair comparisons. In short, we highlight the importance of consistent image resizing for fair evaluation in OOD detection, which we hope will guide future research.
5.4 Near-OOD Tasks
Near-OOD tasks refer to setups where the inlier and outlier samples are semantically similar, making them challenging for most methods. From Table 3, DiffPath, like most baselines, does not perform strongly on near-OOD tasks like CIFAR10 vs CIFAR100. This motivated us to conduct further near-OOD experiments, the results of which are presented in Table 8 of Sec. C of the appendix. Note that near-OOD tasks are not a standard evaluation on generative methods. We defer detailed analysis of the results to the appendix, and leave further investigations on near-OOD tasks to future work.
@@ 6 Conclusion
In this work, we proposed Diffusion Paths (DiffPath), a method of OOD detection using a single diffusion model by characterizing properties of the forward diffusion path. In light of the growing popularity of generative foundation models, our work demonstrates that a single diffusion model can also be applied to OOD detection. There are several interesting future directions that arise from our work; for instance, applying DiffPath to other modalities such as video, audio, language, time series and tabular data, as well as investigating if higher-order terms of the Taylor expansion, or leveraging different instantiations of the PF ODE might lead to better performance.
@@
We only calculate the simple time derivative and found that it works well experimentally, although one might compute the full derivative to quantify the curvature completely. We leave this for future work. Also, we consider DMs trained on natural images like CelebA and ImageNet, which may not be appropriate for specialized applications such as medical images. For such purposes, one could consider including domain-specific data during training.
@@ 7 Acknowledgements
This research is supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-RP-2020-017). AHT acknowledges support from the Singaporean Ministry of Education Grant MOE-000537-01 and MOE-000618-01. We would like to thank Pranav Goyal for help with the experiments.
Supplementary Material for “Out-of-Distribution Detection with a Single Unconditional Diffusion Model”
@@ Appendix B Experimental Details
As mentioned in Sec. 5, we utilize a single unconditional diffusion model trained on CelebA and ImageNet at $32\times 32$ and $64\times 64$ resolution respectively. We train our own CelebA model and utilize the ImageNet checkpoint trained using Improved DDPM’s $L_{\textrm{hybrid}}$ objective (Eq. 16 of Nichol and Dhariwal [40]) from the official repository [^2] https://github.com/openai/improved-diffusion. Both models use a cosine noise schedule with a total of 4000 diffusion steps. For DiffPath-1D, we fit a KDE using a Gaussian kernel with a bandwith of 5. For DiffPath-6D, we fit a GMM with hyperparameters obtained by sweeping over a predefined number of mixture components (e.g., 50, 100) and covariance type (e.g., diagonal, full, tied). Both are implemented using the sklearn library.
On a single Nvidia A5000 GPU, DiffPath takes approximately 0.25s and 0.94s per integration step on $32\times 32$ and $64\times 64$ images respectively with a batch size of 256.
Diffusion Baselines.
@@
For all diffusion baselines, we rely on the official GitHub repositories and open-source checkpoints where possible. The repositories are listed as follows: MSMA [^3] https://github.com/ahsanMah/msma, DDPM-OOD [^4] https://github.com/marksgraham/ddpm-ood, LMD [^5] https://github.com/zhenzhel/lift_map_detect. For NLL and IC, we use the pre-trained CIFAR10 checkpoint from Improved DDPM and train our own model for SVHN using the same hyperparameters at $32\times 32$ resolution. We calculate the NLL using the default implementation in Improved DDPM, while we compute IC using code from the LR repository [^6] https://github.com/XavierXiao/Likelihood-Regret due to lack of official code from the IC authors. We train all baselines using 1-3 A5000 GPUs.
Appendix C Near-OOD Results
Table 8: Average AUROC results for near-OOD tasks as proposed in OpenOOD [41]. We use DiffPath-6D with ImageNet as the base distribution with 10 DDIM steps. Bold and underline denotes the best and second best result respectively. We also show the number of function evaluations (NFE) for diffusion methods, where lower is better.
Method CIFAR10 TinyImageNet
KLM 0.792 0.808
VIM 0.887 0.787
KNN 0.907 0.816
DICE 0.783 0.818
DiffPath-6D 0.607 0.845
To further investigate the performance of DiffPath on near-OOD tasks, we ran experiments on two tasks proposed in OpenOOD [41]. The first task involves CIFAR10 as the in-distribution data with CIFAR100 and TinyImageNet (also known as ImageNet200) as out-of-distribution datasets. The second involves TinyImageNet as in-distribution data and SSB [42] (hard split) and NINCO [43] as out-of-distribution data (see the official repository [^7] https://github.com/Jingkang50/OpenOOD for full details).
We compare our results against the four latest discriminative baselines reported in OpenOOD [41] under the “w/o Extra Data, w/o Training" category, which are KLM [44], VIM [45], KNN [46] and DICE [47]. We report the average AUROC of each task in Table 8. The results are mixed: DiffPath performs the best for the TinyImageNet task, but obtains the poorest result for the CIFAR10 task.
As noted in the main text, to our knowledge near-OOD tasks are not a standard evaluation setup for generative methods. We hypothesize that such tasks are better suited to discriminative methods as gradient-based classification training enables the model to learn fine-grained features specific to each in-distribution class, which we believe is crucial in this context. In contrast, generative models focus on maximizing the likelihood of the overall data distribution and are not explicitly trained to identify subtle discriminative features. This could potentially lead to weaker performance in tasks where the distributions exhibit a high degree of overlap. It should be noted that discriminative methods typically require class labels, while unconditional generative methods do not, thus constraining the use of the former to cases with labelled in-distribution data.